This is an Example of the Title of a Fictitious Report

First Author, Second Author, Third Author

Report A/200N/M

ISBN NNNNNNNNNN ISNN 0787-6416

UNIVERSITY OF KUOPIO Department of Computer Science

P.O.Box 1627, FIN-70211 Kuopio, FINLAND